OpenStudy (anonymous):

What is the volume of a triangular prism with a height of 7 inches and a base that is an equilateral triangle with sides of 6 inches?

OpenStudy (anonymous):

Since the formula for volume V = B*H /3, where B is area of the base and H is height, you get, when you calculate the area of equilateral triangle (B = 6^2 sqrt(3)/4, B=9 * sqrt(3)), V=(9 * sqrt(3) * 7) / 3, V = 21 * sqrt(3). There you have it:)

OpenStudy (anonymous):

ooo, thank you so much <3